OpenStudy (anonymous):

The slope of the line below is -2. Which of the following is the point-slope form of the line? (4,-8)

OpenStudy (anonymous):

is it y - 8 = -2(x + 4) or...

OpenStudy (espex):

Point-slope form of a line is \[y - y_1 = m(x - x_1)\] Does your equation match this form?